Circular No. 9079

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1647 ORIONIS C. Aspin, Institute for Astronomy, University of Hawai’i (UH); C.

Briceno, J. Downes, F. Hernandez, and R. Rojas, Venezuela National Astronomical Observatory (VNAO); and M. Nakano, Oita University, Japan, report the following optical observations from the VNAO 1-m Stock Telescope (in V and I bands) and the UH 2.2-m telescope on Mauna Kea (in r′ and i′ bands). On Sept. 11 UT, the young eruptive variable star V1647 Ori — emerging from conjunction with the sun — was observed to have magnitudes V = 18.00 ± 0.15, I = 14.55 ± 0.03, and V –I = +3.45 ± 0.17. On Sept. 12 UT, V1647 Ori had magnitudes r′ = 17.69 ± 0.05, i′ = 15.85 ± 0.05, and r′–i′ = +1.84 ± 0.07. These values are close to the maximum values reached by V1647 Ori both soon after outburst in February 2004 (cf. IAUCs 8354, 8396, 8600) and during its subsequent brightening in September 2008 (cf. IAUCs 8968, 8969).

COMETS P/2009 QG_{31}, 226P, AND 227P Comet P/2009 QG_{31} (cf. IAUC 9078) has been given the name “La

Sagra” by the Committee on Small Bodies Nomenclature. Comet P/2009 R2 (= P/1783 W1 = P/2003 A1; cf. IAUCs 9072, 9073) has been assigned the permanent number 226P. Comet P/2009 S4 (= P/2004 EW_{38}; cf. IAUC 9077) has been assigned the permanent number 227P.

COMETS C/2009 F10, C/2009 G2–G6 (SOHO) Further to IAUC 9077, additional Kreutz sungrazers have been found

on SOHO website images. All were very diffuse except for C/2009 G2 (which was small and elongated, with a short, faint tail) and C/2009 G3 (which was small and condensed, peaking at mag ∼ 7 in C3 images, though C2 images showed it to have a hint of a tail and a ‘lump’ suggestive of possible fragmenting). Approximate peak magnitudes in C2 images: C/2009 F10, 8–8.5; C/2009 G2, 7; C/2009 G4, 8.5; C/2009 G5 and C/2009 G6, 8.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 F10 Mar. 30.038 0^{h}43.^{m}5 + 3^{o}09′ C2  MU 2009-N29
C/2009 G2 Apr. 2.104 0 52.4 + 3 52 C2  HS 2009-N30
C/2009 G3  2.113 1 04.4 + 3 47 C3/2 BZ 2009-N30
C/2009 G4  4.646 1 03.9 + 5 13 C2  AK 2009-N30
C/2009 G5  6.963 1 12.1 + 6 02 C2  HS 2009-N30
C/2009 G6  10.329 1 24.6 + 7 11 C2  BZ 2009-N30
2009 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green